OpenStudy (anonymous):

x-1/4=x+23 ....It's really confusing because I don't know what to do I always get lost in the problem

jimthompson5910 (jim_thompson5910):

Is the problem \[\Large \frac{x-1}{4}=x+23\] ???

OpenStudy (anonymous):

X-1/4=X+2/3

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

oh ok So it's \[\Large \frac{x-1}{4}=\frac{x+2}{3}\] ???

OpenStudy (anonymous):

So like don't you have to to Multiply The 4 by 4 and Divide?

jimthompson5910 (jim_thompson5910):

If so, then... (x-1)/4 = (x+2)/3 3(x-1) = 4(x+2) 3x-3=4x+8 3x=4x+8+3 3x-4x=8+3 -x=8+3 -x=11 x=11/(-1) x=-11 So the solution is x = -11

jimthompson5910 (jim_thompson5910):

let me know if anything doesn't make sense

OpenStudy (anonymous):

3(x-1) = 4(x+2) when you did 3(x-1) you distribute right?

jimthompson5910 (jim_thompson5910):

on that step, I was cross multiplying

jimthompson5910 (jim_thompson5910):

when I got 3x-3=4x+8, I was distributing

OpenStudy (anonymous):

Yeah that's what I figured

OpenStudy (anonymous):

yea lol my caculater is messed up!

OpenStudy (anonymous):

THANKS

jimthompson5910 (jim_thompson5910):

yw